### v1

# Predictive Communications for Low-Altitude Networks

Abstract
The emergence of dense, mission-driven aerial networks supporting the low-altitude economy presents unique communication challenges, including extreme channel dynamics and severe cross-tier interference. Traditional reactive communication paradigms are ill-suited to these environments, as they fail to leverage the network’s inherent predictability. This paper introduces predictive communication, a novel paradigm transforming network management from reactive adaptation to proactive optimization. The approach is enabled by fusing predictable mission trajectories with stable, large-scale radio environment models (e.g., radio maps). Specifically, we present a hierarchical framework that decomposes the predictive cross-layer resource allocation problem into three layers: strategic (routing), tactical (timing), and operational (power). This structure aligns decision-making timescales with the accuracy levels and ranges of available predictive information. We demonstrate that this foresight-driven framework achieves an order-of-magnitude reduction in cross-tier interference, laying the groundwork for robust and scalable low-altitude communication systems.
I Introduction
The emergence of the low-altitude economy, an ecosystem of services conducted by unmanned aerial vehicles (UAVs) and electric vertical takeoff and landing (eVTOL) aircraft, is set to redefine urban and rural infrastructure [1, 2]. From automated logistics to real-time infrastructure inspection, these services depend on the seamless operation of extensive low-altitude networks. As these networks form a dynamic aerial layer above the terrestrial systems [3, 4], the critical role of a robust and intelligent wireless communication paradigm in low-altitude networks becomes paramount.
Figure 1: The operational concept of the low-altitude network and its implications for predictive communication. (a) Diverse low-altitude operations, such as infrastructure inspection and cargo delivery, are performed by aircraft following trajectories that are predetermined by their mission requirements. (b) For reasons of air traffic safety and operational management, these diverse, mission-driven flight plans are filed with a central control authority, which possesses a priori, network-wide knowledge of all aircraft movements. This centralized oversight of a predictable, time-varying network topology creates the foundational opportunity for predictive communication.
Figure 2: Illustration of an aerial-assisted radio map generation and update process. Multiple UAVs collaborate to sample the radio environment along their trajectories, creating a sparse point cloud of channel data. This aggregated data is then used to generate or update a dense, multi-layered environmental model, enabling a continuously refined understanding of the radio environments.
However, low-altitude networks present a unique hybrid challenge that is not addressed by existing paradigms. They are far more dynamic than conventional terrestrial networks [5, 6], with high-speed, 3D [^1] Throughout this paper, we use the notation xD to denote an x-dimensional space. mobility that induces drastic channel variations. However, they are more constrained and operate in a much more complex environment than satellite networks. While space systems [7, 8] benefit from near-perfect trajectory predictability in an electromagnetically sparse void, low-altitude aircraft face unpredictable wind, dynamic obstacles, and a dense, interference-coupled airspace that invalidates the assumptions of traditional satellite communication design.
The opportunity to master this environment lies in harnessing a powerful, yet underutilized, resource: imperfect but statistically robust predictability. This foresight is derived from two pillars: the a priori knowledge of mission-driven flight plans filed with a central authority [9, 10, 11], as illustrated in Fig. 1, and the quasi-static nature of the large-scale radio environment, which can be captured in models like radio maps [12, 13, 14], as illustrated in Fig. 2. This available foresight allows us to shift from a reactive to a proactive design philosophy fundamentally.
This paper introduces a layered predictive communication framework that is explicitly designed to harness this imperfect foresight. The key to unlocking unprecedented efficiency and control is to hierarchically decompose the network management problem in a way that matches the intrinsic structure of the predictive information itself. Our framework translates the coarse, long-horizon global predictions into strategic directives, while leveraging precise, short-horizon local data for tactical execution and operational efficiency. Our key contributions are made as follows:
  • We establish foundational principles of predictive communication in low-altitude networks, demonstrating the predictability of both trajectory and radio environment, and illustrating how these can be leveraged to forecast future channel states.
  • We analyze the hierarchical nature of predictive uncertainty, identifying a fundamental trade-off between predictive range and accuracy across central, local, and individual network echelons.
  • We propose a layered predictive communication framework, built upon the above hierarchical understanding. This framework decomposes the predictive cross-layer resource allocation problem into three layers: a strategic layer for network-level routing using coarse information, a tactical layer for group-level timing coordination, and an operational layer for precise link-level power allocation. A case study demonstrates the framework’s effectiveness in mitigating more than 10 dB of cross-tier interference.
II Foundations of Predictive Communication: Building Actionable Foresight
The technical feasibility of predictive communication relies on the ability to forecast future channel states. Two key pillars of predictability make this possible in low-altitude networks: the deterministic nature of mission-driven aircraft trajectories and the statistical stability of the large-scale radio environment. This section formally establishes these foundations. We will first introduce each of the two pillars and then detail the synthesis process that combines them to produce a predictive channel forecast, as illustrated in the workflow of Fig. 3.
Figure 3: Generation process of the predictive large-scale CSI, illustrating the information processing flow. The architecture leverages two foundational inputs: (a) a set of 4D mission trajectories (3D space + 1D time) for all aircraft, and (b) a 6D radio environment model that maps the 6D spatial coordinates of a transmitter-receiver pair to a 1D prediction of large-scale channel quality. In the (c) synthesis step, these inputs are fused. For any two aircraft, their independent 4D trajectories are first temporally aligned to form a 7D state descriptor (3D transmitter position, 3D receiver position, and 1D time). At each time step, the 6D spatial component is mapped by the environmental model to a 1D channel quality value. This process transforms the initial trajectory data into two representative predictive views: a link-level channel prediction, which shows the 2D (channel quality vs. time) forecast for a series of single links, and a network-level channel prediction, which displays the evolution of the entire low-altitude network topology.
II-A Pillar 1: Trajectory Predictability
The first pillar of our predictive communication architecture is the set of predetermined 4D trajectories, that is, three spatial coordinates plus time, for every low-altitude aircraft in the network. The availability and utility of this information are underpinned by two practical facts inherent to the low-altitude networks.
  • Mission-driven flight planning. Most low-altitude aircraft are deployed for clearly defined tasks, ensuring their flight paths are well-planned. These plans are filed with a central control authority for air-traffic safety and operational management reasons.
  • Statistical adherence to the flight plan. Although wind, temporary restrictions, or ad-hoc directives may cause brief deviations, aircraft typically re-converge to their planned routes. Over the timescales relevant to resource planning, the realized trajectory aligns, on average, with the pre-filed path.
These principles are illustrated in Fig.1. Various mission-driven tasks are depicted in Fig.1(a), such as cargo aircraft following fixed logistics routes and inspection aircraft conducting repetitive scans over infrastructure. The architectural implication of this operational model is shown in Fig.1(b), where all flight plans in low-altitude networks are filed with and orchestrated by a central control authority. This entity, therefore, possesses network-wide, a priori knowledge of all planned aircraft movements. Owing to the high statistical adherence to these plans, the conceptual architecture in Fig.1(b) can be regarded as both a static plan and a snapshot of the network’s expected real-time evolution.
Taken together, these two facts allow us to treat the complete set of 4D trajectories as a deterministic and statistically robust forecast of future node locations. In our architecture, this information forms the time-evolving skeleton of the network topology. It serves as a primary input, as illustrated in Fig. 3(a), upon which all subsequent channel predictions are built.
II-B Pillar 2: Radio Environment Predictability
The second pillar of our architecture is a predictable model of the radio environment that maps every transmitter-receiver coordinate pair to a statistical description of the large-scale channel. Formally, it is a static function from a 6D spatial input, the 3D coordinates for the transmitter and the receiver, to CSI statistics, including path-loss, shadowing, and etc. Two observations justify this pillar.
  • Quasi-static large-scale channel characteristics. Over minutes or hours, path-loss and shadowing between two fixed points remain effectively constant, even though small-scale fading fluctuates rapidly.
  • Advanced radio-environment models now provide fine-grained, city-scale accuracy. State-of-the-art radio maps generated via ray tracing, measurement-driven interpolation, or high-fidelity digital twins can deliver metre-level resolution across the 3D workspace [12, 13, 14]
An aerial-assisted radio map is a powerful example of such a model, whose dynamic generation and update process is illustrated in Fig. 2. In this paradigm, aircraft act as mobile sensing agents, collecting geo-tagged channel measurements along their trajectories. This sparse dataset is collected and then transformed into a dense, multi-layered map at a central or edge repository using advanced radio-environment models that infer channel characteristics in unsampled locations. Crucially, this mechanism enables the continuous real-time update of the map, allowing it to adapt to environmental changes and reflect the most current state of the radio landscape.
Taken together, these facts allow us to treat the radio environment as a predictable, static information layer. As illustrated in Fig. 3(b), this enables a central controller to query a database for the expected channel quality between any two predicted future locations, providing the second crucial input for our predictive synthesis.
Network Echelon Predictive Range Dominant Sources of Uncertainty Suitable Task Class Example Proactive Task
Central (e.g., Central Control Authority) Long-Term, Network-Level Trajectory Deviations, Environmental Modeling Errors, Channel Randomness Strategic, Network-Wide Planning End-to-End Routing
Local (e.g., Base Station, Cluster Head) Middle-Term, Group-Level Environmental Modeling Errors, Channel Randomness Tactical, Multi-Node Coordination Handover Timing
Individual (e.g., Aircraft) Short-Term, Link-Level Channel Randomness (after measurement) Link-Level Policy Optimization Power/Rate Control
Table I: The Hierarchical Nature of Predictive Information and Corresponding Proactive Tasks.
II-C Synthesis: Fusing Trajectories and Environment model
The synthesis of the two architectural pillars is a computational process that transforms deterministic knowledge about geometry and movement into actionable foresight about future communication quality. As depicted in Fig. 3(c), this is achieved by systematically evaluating the static radio environment model (Pillar 2) at the future coordinates dictated by the time-varying mission trajectories (Pillar 1). This fusion effectively maps the dynamic network skeleton with predicted channel quality values, creating a unified, time-varying forecast of the network’s CSI.
The output of this synthesis provides two complementary and essential predictive views. At a granular level, it yields link-level channel forecasts, which are time-series predictions of the quality for any specific link, as illustrated in the plot in Fig. 3(d). At a holistic level, the aggregation of all such forecasts produces the final architectural output: the predictive, spatio-temporal channel graph (Fig. 3(e)). This is achieved by mapping the predicted channel quality between every pair of nodes at each time instant to the weight of the corresponding edge in a network graph for that moment. This rich, 4D data structure represents the complete, time-indexed evolution of the network’s connectivity and link qualities.
The successful generation of this graph is not merely a technical step; it is the proof of concept that establishes the fundamental feasibility of the predictive paradigm.
III Hierarchical Nature of Predictive Information
The predictive architecture outlined in the II provides an idealized glimpse of future connectivity; however, real-world forecasts are inherently probabilistic. Factors such as modeling errors, trajectory deviations, and feedback delays introduce uncertainty that grows with the prediction horizon. Moreover, the information available to different network entities, from a central controller to an individual aircraft, varies in scope, freshness, and fidelity, yielding a hierarchical information structure governed by a fundamental range-accuracy trade-off. This section identifies the principal sources of uncertainty, presents a three-tier hierarchy of predictive entities, and shows how their distinct sensing and processing capabilities lead to the trade-off summarized in Table I.
III-A Fundamental Sources of Predictive Uncertainty
The gap between an idealized forecast and real-world performance stems from three primary sources of uncertainty, each degrading predictive quality in a different way:
  • Trajectory deviations: Although missions are pre-planned, an aircraft’s realized path can drift from its filed trajectory because of wind gusts, temporary air-traffic directives, or minor navigation corrections. Even small lateral shifts are critical in the low altitude networks: a displacement of only a few meters may flip a link from line-of-sight (LOS) to NLOS, causing significant errors in large-scale channel prediction.
  • Environmental model inaccuracy and staleness: State-of-the-art radio maps and digital twins provide high-fidelity, large-scale channel estimates, yet they are still abstractions of reality. In addition, the global model held by a central controller is updated only as frequently as it receives, often compressed, reports from aircraft. Consequently, its view is less fresh and less accurate than the locally refined model maintained on board each platform.
  • Channel randomness: Even with perfect location knowledge and a flawless environment model, wireless channels exhibit unpredictable small-scale fading induced by multipath propagation. This represents a fundamental physical limit: fine-grained channel states cannot be known without instantaneous measurement.
Together, these factors transform deterministic foresight into probabilistic insight, motivating the layered optimization framework developed in the sequel.
Figure 4: The layered predictive communication framework, illustrating the hierarchical task decomposition for a proactive communication task. (a) The overall objective is to deliver data from a source to a destination via a multi-hop aerial route while minimizing interference. The main decisions, route selection, timing, and power allocation, require information at different scales. This complex problem is decomposed across three layers: (b) The Large-Scale Strategic Layer uses global, long-term information to transform mission requirements into a reserved end-to-end path. (c) This path is passed as a directive to the Middle-Scale Tactical Layer, where a local cluster uses more precise, local-area predictions to coordinate the timing of each hop, potentially adjusting the route to bypass a blockage. (d) Finally, the Small-Scale Operational Layer receives a specific transmission schedule. It uses link-level statistics to design a predictive power allocation policy that optimizes physical-layer efficiency while managing interference.
III-B Accuracy vs. Range Trade-Off
The interplay between the physical capabilities of the network echelons and the fundamental sources of uncertainty gives rise to a distinct information hierarchy, defined by a crucial trade-off between predictive accuracy and range.
The central echelon is characterized by the maximum predictive range but the lowest accuracy. At the apex, the central echelon, typically a central control authority, retains every pre-filed 4D trajectory and a city-scale radio map, offering a network-wide, long-horizon predictive view. Its range is a direct consequence of its physical role as the global repository for all pre-planned mission trajectories, which grants it an unparalleled, long-horizon view of the entire network’s planned evolution. However, this reliance on a global, planned-level abstraction is the source of its low predictive fidelity. Its model must operate on planned rather than real-time trajectories, forcing it to treat on-the-fly deviations as a statistical variance. Furthermore, its large-area environmental model is subject to information staleness and inaccuracies from periodic, aggregated reporting. Consequently, the central forecast must account for the full spectrum of these uncertainties, yielding a high-variance statistical model.
The local echelon, in contrast, exhibits intermediate range and significantly improved accuracy. Serving as an intermediary, the local echelon, often a ground base station or edge server, aggregates high-rate feedback, such as the real-time positions of neighboring nodes, from a regional cluster of aircraft, refining the central forecast with low-latency, high-fidelity local data. Its range is confined to a local subgraph, such as the aircraft within a base station’s coverage area. Its physical capability as a regional information aggregator allows it to gather real-time positional data from its local cluster, thereby eliminating the trajectory deviation uncertainty that plagues the central tier. While its predictive range is limited to its regional sub-graph, the resulting lower-variance forecast is more reliable.
Finally, the individual echelon has the most limited range but the highest accuracy. At the base is the individual echelon, consisting of each aircraft itself; equipped with onboard sensors, it performs instantaneous channel measurements and precise self-localization. Its predictive scope is restricted to its egocentric, one-hop neighborhood. However, its onboard capability to perform direct, instantaneous channel measurements provides a ground-truth anchor for its forecasts. This allows it to temporarily eliminate all sources of uncertainty, yielding a predictive model with a unique, near-zero initial variance.
Therefore, this clear hierarchy of capabilities logically dictates a corresponding function hierarchy. The central tier’s broad but coarse foresight is well-suited for strategic, network-wide planning. The local tier’s balance of regional scope and tactical-grade fidelity is ideal for multi-node tactical coordination. Lastly, the individual tier’s highly precise but myopic foresight perfectly matches the link-level policy optimization task. This functional decomposition, summarized in Table I, forms the guiding principle for our proactive control framework.
Figure 5: Illustration of large-timescale proactive interference-aware path reservation. (a) The considered scenario, showing aircraft mission trajectories with directions (red arrows) within a ground layout that includes a source (square), a destination (triangle), and interference-sensitive nodes (circles). (b) The computed route and associated transmit power (colorbar) for a delay-tolerant task (delay tolerance is 20s). (c) The computed route and transmit power for a delay-sensitive task (delay tolerance is 2s). (d) shows the aggregate interference power of our proposed method versus two baselines under increasing network load.
IV Layered Predictive Communication Framework
The analysis in the accuracy-range trade-off dictates a new design philosophy: one that embraces the hierarchical nature of predictive information as a structural blueprint for the optimization framework. This leads directly to our proposed solution: a layered predictive communication framework. The core principle of this framework is hierarchical task decomposition, as conceptually illustrated in Fig. 4. The predictive cross-layer objective of delivering data with minimal interference (Fig.  4(a)) is broken down into a cascade of distinct, tractable sub-problems. This is achieved by assigning different functional responsibilities to three layers, each operating at a different scale and leveraging the predictive information best suited to its task. This hierarchical decomposition is the key to creating a system that is both strategically robust and tactically agile.
IV-A The Large-Scale Strategic Layer
The strategic layer, operating at the apex of the hierarchy, performs long-term, network-wide path reservation. This centralized function is critical, as only this layer possesses the global, long-horizon view necessary to optimize end-to-end routes, a task beyond the localized scope of lower echelons. While its predictive model is statistically coarse, it effectively establishes robust, high-level directives. The architecture is designed to dynamically refine this initial imprecision through the more agile tactical and reactive layers below. As illustrated in Fig. 4(b), the strategic layer translates global mission requirements into reserved paths using the global spatio-temporal channel graph, which serves as a foundational strategic directive for the entire network.
IV-B The Middle-Scale Tactical Layer
Operating under the directives of the strategic layer, the tactical layer is responsible for coordinating multiple nodes within a localized context. This layer leverages more accurate, mid-term, local-area predictions to translate the high-level strategic plan into a refined set of concrete, coordinated actions. Its key function is to adapt the strategic plan to the dynamic realities of the local environment. A canonical example is illustrated in Fig. 4(c). The strategic layer may have issued a route directive of $2\to 6\to 7$. However, the nodes in the local cluster observe that the direct link $2\to 6$ is currently blocked or experiencing deep fading. Therefore, the tactical layer’s first action is to perform a local re-routing, adjusting the path to $2\to 5\to 7$ to bypass the obstruction. Having established a viable new path, it then proceeds to coordinate the transmission timing. Using their superior local predictive models, nodes 2 and 5 can identify and schedule their respective transmissions to coincide with periods of high channel gain. This dual capability of reactive re-routing and proactive re-timing allows the system to enhance efficiency and resilience in a way the coarse global model could not foresee.
IV-C The Small-Scale Operational Layer
At the base of the hierarchy, the operational layer tackles short-term, link-level optimization. Each aircraft, equipped with on-board sensors, uses the most accurate measurement-anchored statistics to pre-compute a mapping from instantaneous channel states to physical-layer parameters. In Fig. 4(d), the node leverages high-fidelity predictions of its desired link and the air-to-ground interference link to derive an interference-aware power control rule ahead of transmission. By exploiting precise channel information, the operational layer enables low-latency, energy-efficient, and interference-aware resource allocation.
This layered decomposition is the key to creating a strategically robust and tactically agile system. The layers operate in a top-down cascade of goal-setting and refinement: the strategic layer’s global routing plan defines the high-level objectives and constraints for the tactical layer below. In turn, the tactical layer’s coordinated schedule provides a specific, timed mission for the operational layer. This structure enables the network to develop robust, long-term plans based on statistical foresight, while maintaining the flexibility to execute those plans with precision in response to real-world conditions. The practical efficacy of this hierarchical approach is validated in the subsequent case study.
V Case Study: Interference-Aware Predictive Communication
This section validates our layered proactive framework through a comprehensive case study focused on predictive interference mitigation [15], with key results illustrated in Fig. 5. First, the large-scale layer demonstrates strategic adaptability, exploiting aircraft mobility for delay-tolerant tasks (Fig. 5(b)) while using multi-hop relays for delay-sensitive tasks (Fig. 5(c)). The middle-scale layer then performs tactical timing coordination based on channel dynamics, scheduling a high-power, rapid transmission for the degrading first two hops in Fig. 5(b) while using a low-power, longer transmission for the improving last hop. The small-scale layer ensures fine-grained interference control, strategically reducing transmit power when aircraft operate near sensitive ground nodes, a behavior confirmed by the colorbar. As validated in Fig. 5(d), this holistic, proactive strategy consistently achieves a significant reduction in aggregate interference power compared to traditional terrestrial (such as aggregate-based communication) and space (such as space-time-based communication) communication baselines, with this advantage becoming more pronounced as network load increases.
VI Challenges and Opportunities
There are multiple future directions for predictive communications in low-altitude networks worthy of further investigation, as discussed below.
VI-A Dynamic Environmental Awareness: Real-Time and Semantic Model Updates
Our framework demonstrates the power of leveraging static environmental models. A critical next step is to enable these environmental representations to dynamically update in real time, transforming them into "living" models of the world. This introduces key challenges in efficient data management. Semantic communication techniques are needed on the uplink to allow aircraft to intelligently identify and report only the most salient environmental changes (e.g., terrain and weather), rather than raw CSI measurements. On the downlink, efficient dissemination of model updates is crucial, potentially by partitioning the model into standard (global) and private (local) components to reduce redundancy. This would create a system with continuously evolving environmental awareness, making its predictions even more accurate and adaptive.
VI-B Robust Decentralized Execution
While our framework provides a strategic plan, its execution can be made more robust against unforeseen events, such as node failures or unexpected blockages. A vital future direction is to enhance system resilience by moving towards decentralized execution models, likely using multi-agent reinforcement learning. The core research challenge is to develop decentralized policies that enable aircraft to follow the central plan and autonomously re-coordinate to find alternative solutions when disruptions occur, ensuring mission continuity. Techniques like knowledge distillation could be explored to accelerate this adaptive learning process.
VI-C Predictive Coverage
The predictive paradigm can be extended beyond routing and resource allocation to optimize network coverage itself proactively. This can be approached from two angles. First, predictive interference maps can inform the long-term planning and adjustment of terrestrial base station parameters (e.g., antenna tilts) to serve heavy-traffic air corridors better. Second, the framework can guide the dynamic deployment and trajectory planning of aerial base stations, positioning them not where coverage is currently weak, but where it is predicted to be needed, proactively filling anticipated gaps in terrestrial service.
VI-D Predictive integrated sensing and communication (ISAC)
Integrating sensing into the predictive framework offers a powerful way to mitigate the massive data redundancy inherent in multi-aircraft ISAC. The key is to co-schedule sensing and communication tasks by leveraging foresight in both temporal and spatial dimensions. Temporally, the framework can predict future windows of high-quality sensing channels (e.g., optimal target angle) and communication channels, enabling a sparse schedule of high-value sensing actions and opportunistic, low-power transmissions for a single aircraft. Spatially, it can predict future overlaps in the sensing fields of view over multiple aircraft. This foresight enables proactive coordination, such as assigning specific sensing sectors to each aircraft or selecting a single node to measure. In contrast, others relay the data, thus eliminating redundant sensing and transmission.
VII Conclusion
This paper pioneered a predictive communication paradigm for low-altitude networks, proposing a novel layered framework that translates hierarchical predictive information into robust strategies for routing, coordination, and policy design. To spur further innovation, several key future research directions were also identified to support and promote the continued development of this predictive framework.